OpenStudy (anonymous):

Find the Derivative of csc2x/e to negative x

OpenStudy (anonymous):

Is the question: \[\frac{d}{dx}\frac{\csc(2x)}{e^{-x}}\]

OpenStudy (anonymous):

sorry for the late reply and yes it is

OpenStudy (anonymous):

Okay, well we can simplify as: \[\frac{d}{dx}\frac{\csc(2x)}{e^{-x}}=\frac{d}{dx}\frac{1}{\sin(2x)e^{-x}}=\frac{d}{dx}[\sin(2x)e^{-x}]^{-1}=\frac{d}{dx}[f(x)g(x)]^{-1}\] Now, we can evaluate this as: \[\eqalign{ &=-[f(x)g(x)]^{-2}[f(x)g(x)]' \\ &=-\frac{1}{[f(x)g(x)]^2}\times[f(x)g'(x)+g'(x)f(x)] \\ &=-\frac{f(x)g'(x)+g'(x)f(x)}{f(x)^2g(x)^2} }\]

OpenStudy (anonymous):

I got e